Question 6 12 pts An unknown reaction has a rate constant value of 0.202 s 1 at 228 °C, and of 1.129 s1 at 367 °C. What is th   can you solve and show work if possible
Answer #1

k, = 0.2025! T= 2286 -7 (228 + 243-15) K T, = 501-15K - K2= 1.129 51 Tz = 3676 =) (367 +273.15) K 640:15K let activation ener